The Earth’s orbit is of mean radius 1.50 x 1011 m and the Earth’s year is 365 days long.

(i)The mean radius of the orbit of Mercury is 5.97 x 1010 m. Calculate the length of Mercury’s year.

(ii) Neptune orbits the Sun once every 165 Earth year.

    Calculate the ratio: distance from Sun to Neptune / distance from Sun to Earth

1
Expert's answer
2021-02-25T11:29:20-0500

Answers

Time period of mercury can be given as

"T_m=\sqrt{\frac{R_m^3T_e^2}{R_e^3}}"

"T_m=\sqrt{\frac{(5.97\times10^{10})^3(365^2)}{(1.50\times10^{11})^3}}"

"T_m=23days \quad of\quad earth"